A reaction is performed to study the oxidation of nitrogen monoxide by oxygen at 25 °C:

  • 2 NO + O2ebab9d75-6e9a-4c80-9bcd-b491b668f097.gif2 NO2

The following reaction rate data was obtained in four separate experiments.

Experiment [NO]0, M [O2]0, M Initial Rate, M/s
1 9.10×10-3 4.40×10-3 3.11×10-3
2 1.82×10-2 4.40×10-3 1.24×10-2
3 9.10×10-3 8.80×10-3 6.22×10-3
4 1.82×10-2 8.80×10-3 2.49×10-2



What is the rate law for the reaction and what is the numerical value of k?

Complete the rate law in the box below.
Remember that an exponent of '1' is not shown and concentrations taken to the zero power do not appear.

Rate =



k = M-2s-1
Answer #1

2 Nota - 2NQ 9=K[NO] [Q] 3-11 * 103 = K ( 9-1X102) (4:44153) 70 124x10²=k (1.82x102) M (4.4 x103) 1. 6. 22X103 = K (9.1x103